# TuRTLe: A Unified Evaluation of LLMs for RTL Generation

### v1

@@ II-D Integrated Benchmarks
TuRTLe initially integrates a selection of four benchmarks, selected by quality, variety and size. These target specific tasks and align with different design goals introduced in our work. Table II provides a comparative overview of the design goals covered by each integrated benchmarks, together with the number of problem descriptions included. This number is reported separately for LCA (left) and the rest of goals (STX, FNC, SYN and PSQ, right), as they have different requirements.
@@
This benchmark consists of 156 problems sourced from HDLBits [^2] https://hdlbits.01xz.net/wiki/Problem_sets, designed to assess LLMs’ performance in two tasks: MDC and S2R. In the MDC task, models are provided with a problem statement along with a module header and are required to generate the missing body of the module. In the S2R task, models receive a prompt resembling a high-level design specification and must generate an entire module from scratch. VerilogEval was originally designed to evaluate STX, FNC and SYN design goals. TuRTLe extends its evaluation scope by incorporating PSQ. We manually corrected the reference implementation of six problems, to address synthesizability issues as explained in Appendix IV.
@@
This benchmark introduces 50 human-crafted designs of varying complexity, designed to evaluate LLMs on S2R. Originally, RTLLM was developed to assess LLM performance across three evaluation goals: syntax correctness (our SYN), functional correctness (our FNC), and design quality including power, performance, and area metrics (our PSQ). Notice RTLLM definition of syntax correctness is differently from the STX goal used in this work. Specifically, RTLLM evaluates syntax correctness based on whether a design can be successfully synthesized into a netlist without syntax errors which is effectively combining aspects of both STX and SYN as defined in this work. RTLLM originally employs a “success rate metric,” which is conceptually similar to the Pass@$k$ metric [38] used in this work, but it does not strictly adhere to the requirement that $N>k$. We corrected the reference implementation of four problems and excluded two to address functional and synthesizability issues as explained in Appendix IV.
@@ III Experiments
This section presents a comprehensive evaluation of LLMs for RTL generation, analyzing their performance across multiple tasks and design goals using the TuRTLe framework. The following subsections detail the experimental setup, benchmarking methodology, and key findings from our evaluation.
III-A Experimental Setup
TuRTLe evaluates a diverse set of open LLMs, categorized into three groups based on their specialization. The first category comprises general-purpose LLMs, such as LLaMA [41] and DeepSeek [42], which are designed for broad language understanding but lack specific optimizations for code generation. The second category includes coding-specific LLMs, such as QwenCoder [43] and OpenCoder [44], which are fine-tuned on programming tasks. Finally, the third category consists of RTL-specific LLMs, such as CodeV [45] and HaVen [46], which are explicitly optimized to generate Verilog and other RTL constructs more accurately. By evaluating models from all three categories, this study provides a comprehensive understanding of the ability of different type of LLMs for specific RTL generation tasks.
@@ III-A Experimental Setup
We use the Pass@$1$ metric to measure the probability that a model generates a correct solution on the first attempt. However, its accuracy is dependent on the number of samples ($N$) used per evaluation. To determine the optimal value of $N$, we performed ten independent runs on five selected models using varying sample sizes of $N={1,3,5,10,20}$. As shown in Figure 1, the variance decreases as $N$ increases, confirming that small sample sizes lead to unstable results. While larger $N$ values further reduce variance, they impose significant computational overhead. A sample size of $N=5$ is chosen as the optimal configuration, balancing computational efficiency with statistical stability.
@@
Figure 1: Pass@$1$ variance among ten runs while increasing sample size $N$.
TABLE III: Pass@1 performance (higher is better) with n=5 for line and module completion (top table) and specification-to-RTL (bottom table). Each table is split vertically in three, including general purpose LLMs (top), coding LLMs (middle) and RTL-specific LLMs (bottom).
Line Syntax Functionality Synthesis Post-Synthesis Quality
Models RTL-Repo VerilogEval (MC) VeriGen VerilogEval (MC) VeriGen VerilogEval (MC) VeriGen VerilogEval (MC) VeriGen
Power Perf. Area Power Perf. Area
DeepSeek R1 33.02 97.95 94.12 80.26 60.00 79.62 54.12 39.35 38.16 39.21 26.62 26.97 28.01
Llama 3.1 405B FP8 33.29 91.41 72.94 44.74 45.88 44.10 44.71 21.98 20.74 21.66 19.24 22.19 21.08
Qwen2.5 72B 37.19 81.67 70.59 53.08 27.06 52.56 27.06 26.08 24.92 25.83 12.74 13.50 13.89
StarChat2 15B v0.1 13.24 81.54 92.94 39.36 50.59 38.59 50.59 19.21 18.28 19.05 24.02 25.23 25.73
[…]
QwenCoder 2.5 32B 30.44 84.87 72.94 45.51 41.18 44.87 41.18 22.26 21.48 22.20 20.56 20.67 20.87
DeepSeek Coder 33B 30.58 78.72 83.53 39.49 29.41 38.33 29.41 18.92 18.20 18.76 14.52 14.74 14.67
OpenCoder 8B 16.63 79.87 92.94 36.03 43.53 35.51 37.65 17.57 16.74 17.52 17.19 18.76 19.06
[…]
OriGen DeepSeek Coder 19.45 79.36 87.06 43.08 35.29 42.95 35.29 21.50 20.13 21.33 16.55 17.70 18.35
HaVen-CodeQwen 25.38 93.33 97.65 50.00 48.24 48.72 42.35 23.37 23.39 23.09 20.21 21.15 21.25
CodeV-CL-7B 12.39 91.92 98.82 36.79 44.71 36.41 38.82 18.15 16.88 18.05 19.06 19.38 19.35
CodeV-QW-7B 20.56 93.85 57.65 52.56 25.88 51.15 20.00 25.64 24.22 25.56 9.39 9.99 9.94
Syntax Functionality Synthesis Post-Synthesis Quality
Models VerilogEval (S2R) RTLLM v2.0 VerilogEval (S2R) RTLLM v2.0 VerilogEval (S2R) RTLLM v2.0 VerilogEval (S2R) RTLLM v2.0
Power Perf. Area Power Perf. Area
DeepSeek R1 96.54 91.43 79.74 67.76 78.97 63.27 38.94 37.82 38.76 35.64 31.95 34.50
Llama 3.1 405B Instr. 89.10 65.71 57.05 37.55 56.67 35.92 27.18 27.0 26.79 19.70 16.04 18.91
Qwen2.5 72B 81.15 82.04 51.15 47.35 50.38 46.53 25.40 23.83 24.52 24.83 23.88 25.46
StarChat2 15B v0.1 86.54 85.71 38.72 42.45 38.59 42.45 19.18 17.99 19.00 22.44 21.03 22.53
[…]
QwenCoder 2.5 32B 87.69 79.59 45.64 43.27 43.33 42.04 21.51 20.72 21.17 22.02 20.95 22.03
DeepSeek Coder 33B 57.82 83.67 19.87 43.67 19.87 42.86 9.94 9.83 9.47 23.28 21.19 23.20
OpenCoder 8B 75.77 75.10 28.59 46.53 28.21 42.86 13.81 13.16 13.71 22.24 21.47 21.73
[…]
OriGen DeepSeek Coder 90.26 23.67 46.54 12.65 46.92 10.61 23.38 22.18 23.44 5.33 4.61 4.79
HaVen-CodeQwen 90.26 82.45 45.90 40.41 44.36 38.37 21.77 21.23 21.46 19.10 18.31 18.92
CodeV-CL-7B 55.38 69.80 27.05 37.14 26.79 35.10 13.20 12.39 13.03 18.92 16.88 17.89
CodeV-QW-7B 41.79 71.02 19.10 35.51 18.72 27.76 9.36 9.36 9.38 14.85 12.21 13.78
Figure 2: Average STX (left) and FNC (right) scores on VerilogEval and RTLLM across three model categories (General, Coding, and RTL-Specific) at different temperature settings (0.2, 0.5, 0.8). Darker shades indicate higher scores.
III-B Benchmarking Insights
@@
We examine the impact of stochasticity on LLM performance in RTL generation. Temperature controls the randomness of model-generated outputs, influencing both creativity and precision. To assess its effect, we evaluate model performance across three temperature settings: 0.2, 0.5, and 0.8 for Pass@$1$, focusing on STX and FNC goals. Figure 2 shows temperature = 0.2 yields the highest accuracy across all model categories. This suggests that minimal randomness is optimal for RTL generation tasks. Consequently, 0.2 is set as the default temperature for all subsequent experiments to ensure consistency and reproducibility.
Impact of the Context in the Prompt
We analyze how the length of the context from the Verilog project that is given in the prompt affects performance in the single line completion task. We do this by evaluating six models, spanning general-purpose, coding and RTL-specific, using context sizes of 2,048, 4,096, and 8,192 tokens. Our results show that longer contexts consistently improve performance, confirming that LLMs benefit from additional context length when making predictions. Increasing the context length from 2,048 to 4,096 tokens results in a notable improvement of +3.35 in Exact Match scores with a standard deviation of 1.02. Extending the context further to 8,192 tokens provides an additional boost of +2.40 (standard deviation of 1.16). Based on these findings, 8,192 tokens is selected as the standard context length for the SLC task while evaluation RTL-Repo.
@@
VerilogEval supports both MC and S2R tasks using the same underlying dataset, allowing for a direct comparison of how problem definition and presentation impact model performance. Figure 3 compares the performance of base and instruct-tuned variants across ten models spanning five distinct model families: OpenCoder [44], Qwen 2.5 [47], QwenCoder 2.5 [43], DeepSeek Coder V1 [42], and Llama 3.1 [48]. Our results show that there is a clear trend: base models tend to perform better in MC task, where they leverage learned syntax patterns to complete partial module definitions, whereas instruct-tuned models excel in S2R tasks, benefiting from their training on instruction-following tasks that improve their ability to interpret high-level specifications. Given computational efficiency and space constraints, we report base model performance on MC and SLC task, while instruct-tuned models are evaluated on S2R tasks, ensuring that each variant is assessed in the most relevant scenario for its capabilities.
III-C Model Results
This section presents a detailed evaluation of LLM performance across the five key design goals. Table III-A provides a structured comparison, with the top section reports the performance of various models on LCA and MC tasks, while the bottom section highlights results for the S2R task.
Model Performance across RTL Tasks
DeepSeek R1 emerges as the best-performing model across most evaluation metrics, largely due to its autoregressive reasoning chain, which enables iterative refinement of generated code. This advantage is particularly evident in tasks requiring step-by-step reasoning, such as MC and S2R, where DeepSeek R1 significantly outperforms other open models. However, this improvement comes at the cost of increased computational overhead during inference. On RTLLM, for instance, DeepSeek R1 required an average of 5,606 additional tokens solely for its reasoning chain, leading to substantially longer generations. To accommodate these demands, DeepSeek R1 was evaluated using a context length of 16,384 tokens, whereas other models performed effectively with 2,048 tokens. On the other hand, in scenarios where reasoning chains are not employed, such as in real-time SLC task, DeepSeek R1 performs comparably to other top-performing models, suggesting that its primary advantage lies in multi-step reasoning rather than immediate completion accuracy.
A clear distinction is observed in model specialization. General-purpose models perform well in SLC and S2R tasks, where broad contextual understanding plays a crucial role. On the other hand, RTL-specific models excel in MC tasks, leveraging their domain-specific training to generate functionally accurate Verilog modules. Coding-specific models, such as QwenCoder 2.5 32B, demonstrate stable performance across multiple tasks, balancing generalization with code-optimization capabilities.
Figure 4: Average model performance across design goals for specific-to-RTL task.
@@ III-C Model Results
Figure 4 shows average model performance across four design goals (STX, FNC, SYN, and PSQ) for specific-to-RTL task evaluated across all benchmarks. A clear performance degradation is observed as the evaluation criteria become progressively challenging. On average, models achieve a relatively high STX score of 73.88%, indicating that they generate syntactically correct Verilog in most cases. However, transitioning from syntax correctness to functional correctness results in a substantial drop, with the FNC score averaging 40.11% which is a sharp decrease of 33.76%. This decline suggests that while models are proficient at producing Verilog code that adheres to syntax rules, they struggle to generate functionally valid implementations. The significant gap between STX and FNC highlights a key limitation of current LLMs.
Regarding PPA, the PPA-score presented in §II-C has the characteristic that each model has a different human baseline, equal to their SYN score divided by 2. This is because all designs that do not pass synthesis are considered as failures, which ensures the consistency of the whole framework (every step is equal or lower than its predecessor), and illustrates the current limitations of LLMs and benchmarks assessing them. Having this into account, it can be seen in Table III-A that all models perform very close to their human baseline. This could be caused by the simplicity in the synthesized designs, as there is limited room for either significant optimization or flagrant mistakes in simple examples. However, this view is still useful as a sanity check, ensuring that the code produced is at least as optimal as the human reference.
Figure 5: Spearman correlation among benchmarks for SYX and FNC goals.
Benchmark Correlation Analysis
Figure 5 presents the Spearman correlation between benchmarks for STX and FNC, which measures the monotonic relationship between rankings, helping to determine how similarly models perform across benchmarks. The highest correlation observed is between VerilogEval S2R and VeriGen, a moderate positive correlation (0.49). The rest of correlations range between that value and the lack of correlation (zero). These results suggest that each benchmark captures different aspects of RTL generation challenges, emphasizing the importance of combining them for obtaining a comprehensive evaluation of LLM-generated Verilog code.
@@ IV Conclusions
This work presents a comprehensive evaluation of open LLMs for RTL generation, assessing their capabilities across various tasks (single-line completion, module completion, and specification-to-RTL) and design goals (line-level contextual accuracy, syntax correctness, functional correctness, synthesizability, and post-synthesis quality). To achieve this, we developed TuRTLe, a unified evaluation framework that integrates multiple benchmarks into a single, automated infrastructure. This framework simplifies experimentation, enables systematic model evaluation, and allows for future extensibility, making it adaptable to new benchmarks and models.
Our results provide key insights into LLM performance for RTL generation. Reasoning-based models, like DeepSeek R1, generally outperform others but incur higher token generation and inference latency. Model structure also influences task performance, as base models excel in module completion, while instruct-tuned models are more effective in specification-to-RTL tasks, benefiting from structured prompt training. Despite achieving high STX, LLMs struggle with FNC design goals, as evidenced by a significant performance drop between these metrics. Many models generate compilable but functionally incorrect Verilog, highlighting a key limitation. To further assess synthesized designs, we introduce a new PSQ metric, which evaluates performance, power, and area, enabling a direct comparison between LLM-generated and human-designed RTL implementations.
Overall, this analysis represents a significant step toward a comprehensive evaluation of LLMs for EDA workflows, providing insights into their capabilities and limitations in RTL generation. However, our findings also underscore the need for more sophisticated and realistic benchmarks that better reflect real-world design challenges, including multi-module architectures and complex module interconnections. The TuRTLe framework source code is released on GitHub [^3] https://github.com/HPAI-BSC/TuRTLe, ensuring reproducibility. A complete leaderboard, including more models and metrics, is hosted in HuggingFace [^4] https://huggingface.co/spaces/HPAI-BSC/TuRTLe-Leaderboard. Moving forward, we plan to expand our framework with additional benchmarks, further refining the evaluation of LLMs in hardware design automation.

### v2

@@ II-D Integrated Benchmarks
TuRTLe initially integrates a selection of four benchmarks, chosen by quality, variety and size. Table II provides a comparative overview of the design goals covered by each integrated benchmarks, together with the number of problem descriptions included. This number is reported separately for LCA (left) and the rest of goals (STX, FNC, SYN and PSQ, right), as they have different requirements.
@@
This benchmark consists of 156 problems sourced from HDLBits [^2] https://hdlbits.01xz.net/wiki/Problem_sets, designed to assess LLM’s performance in two tasks: module completion (MC) and specification-to-RTL (S2R). In the MC task, models are provided with a problem statement along with a module header and are required to generate the missing body of the module. In the S2R task, models receive a prompt resembling a high-level design specification and must generate an entire module from scratch. VerilogEval was originally designed to evaluate STX, FNC and SYN design goals. TuRTLe extends its evaluation scope by incorporating PSQ. We manually corrected the reference implementation of six problems, to address synthesizability issues as explained in Appendix TuRTLe: A Unified Evaluation of LLMs for RTL Generation.
@@
This benchmark introduces 50 human-crafted designs of varying complexity, designed to evaluate LLMs on S2R. Originally, RTLLM was developed to assess LLM performance across three evaluation goals: syntax correctness (our SYN), functional correctness (our FNC), and design quality including power, performance, and area metrics (our PSQ). Notice RTLLM definition of syntax correctness is differently from the STX goal used in this work. Specifically, RTLLM evaluates syntax correctness based on whether a design can be successfully synthesized into a netlist without syntax errors which is effectively combining aspects of both STX and SYN as defined in this work. RTLLM originally employs a “success rate metric,” which is conceptually similar to the Pass@$k$ metric [38] used in this work, but it does not strictly adhere to the requirement that $N>k$. We corrected the reference implementation of four problems and excluded two to address functional and synthesizability issues as explained in Appendix TuRTLe: A Unified Evaluation of LLMs for RTL Generation.
@@ III Experiments
This section presents the comprehensive evaluation of LLMs for RTL generation, analyzing their performance across multiple tasks and design goals using the TuRTLe framework. The following subsections detail the key findings and complete results. Models used are chosen to guarantee a diverse set of open LLMs, belonging to a wide range of categories. This includes general-purpose LLMs, such as LLaMA [40], code-specific LLMs, such as OpenCoder [41], and RTL-specific LLMs, such as CodeV [42]. In addition, we explore base models and instruct models, when both versions are available, such as Qwen [43]. Finally, we consider the most recent family of models, trained to perform autoregressive reasoning chains before producing a final response commonly known as reasoning models, such as DeepSeek R1 [44] [^3] Reasoning models are not tested for SLC, as their long reasoning chains are not adequate for a task that requires fast responses. Only Qwen3-A22B [45] SLC score is computed, as it has been fine-tuned to support both reasoning and non-reasoning modes.. In total, our comparison includes forty LLMs, all executed under identical and reproducible conditions. Closed models are not included in this study, as these conditions could not be guaranteed.
III-A Benchmarking Insights
Pass@$1$ variance
@@ III-A Benchmarking Insights
We use the Pass@$1$ metric to measure the probability that a model generates a correct solution on the first attempt. However, the consistency of such metric is dependent on the number of samples ($N$) used per evaluation. To determine the optimal value of $N$, we performed ten independent runs on five selected models using varying sample sizes of $N={1,3,5,10,20}$. As shown in Figure 1, the variance decreases as $N$ increases, confirming that small sample sizes lead to unstable results. A sample size of $N=5$ is chosen as the optimal configuration, balancing computational efficiency with statistical stability.
@@
Figure 1: Preliminary study conducted on a subset of models, showing the Pass@$1$ variance among ten runs while increasing sample size $N$.
Figure 2: Pass@1 performance (higher is better) with n=5 for the five goals proposed. SLC: Single-Line Completion. MC: Module Completion. S2R: Specification-to-RTL. Table is split vertically in general purpose (top), coding (middle) and RTL-specific LLMs (bottom). Models marked with an asterisk (*) refer to the base version for SLC and MC tasks, and to the instruct version for S2R. The number in parentheses next to the model names indicates the model size, in billions of parameters. Reasoning models appear in italics. Reasoning models are not tested for SLC, except for Qwen3 235B A22B, which also supports non-reasoning mode.
@@
Prior to the main experiments, we examine the impact of stochasticity on LLM performance in RTL generation. Temperature controls the randomness of model-generated outputs, influencing both variance and precision. To assess its effect, we evaluated model performance across three temperature settings: 0.2, 0.5, and 0.8 for Pass@$1$, focusing on STX and FNC goals, and affected by the choice of $N$ in the previous experiment in III-A, we subsequently found that a temperature of 0.2 yielded the highest accuracy across all model categories (see Appendix -C). This suggests that minimal randomness is optimal for RTL generation tasks. Consequently, 0.2 is set as the default temperature for all subsequent experiments to ensure consistency and reproducibility.
Single-Line Completion Context
We analyze how the length of the context given in the prompt affects performance in single line completion tasks. We do this by evaluating six models, spanning general-purpose, coding and RTL-specific, using context sizes of 2,048, 4,096, and 8,192 tokens. Our results show that longer contexts consistently improve performance, confirming that LLMs benefit from additional context length when making predictions. Increasing the context length from 2,048 to 4,096 tokens results in a notable improvement of +3.35 in Exact Match scores with a standard deviation of 1.02. Extending the context further to 8,192 tokens provides an additional boost of +2.40 (standard deviation of 1.16). Based on these findings, 8,192 tokens is selected as the standard context length for the SLC task.
@@
VerilogEval supports both MC and S2R tasks using the same underlying dataset, allowing for a direct comparison of how problem definition and presentation (e.g., either a fill-in-the-blanks or an instruct-following task) impact model performance. This study is of special relevance for comparing base models against their instruct-tuned counterparts, and can help to optimize the selection and benchmarking of LLMs.
@@
For those models where both versions are available, Figure 3 compares the performance of base and instruct-tuned variants. That includes twelve models spanning six distinct model families: SeedCoder [46], OpenCoder [41], Qwen 2.5 [43], QwenCoder 2.5 [47], DeepSeek Coder V1 [48], Llama 3 [49] and CodeLlama [9]. Results show a clear trend: base models tend to perform better in MC task, where they leverage learned syntax patterns to complete partial module definitions, whereas instruct-tuned models excel in S2R tasks, benefiting from their training on instruction-following tasks that improve their ability to interpret high-level specifications and follow them. The only notable exception to this comes from the deepseek-coder-33b instruct model, which performs poorly on the S2R task. Further details on this experiment can be seen in Appendix -D.
To maximize computational efficiency, for those models where both the base and the instruct version are available, we report base model performance on MC and SLC tasks, and the instruct-tuned models performance on S2R tasks.
III-B Model Results
This section discusses the performance results of LLM across the five design goals. See Figure 2 for the complete results, split by general purpose (top), coder LLMs (middle) and RTL-specific models (bottom).
Model Families
Models with include the autoregressive reasoning chain (DeepSeek R1-0528[44], DeepSeek R1 [44], Qwen3 235B A22B[45], QwQ 32B[50], DeepSeek R1 Distill Qwen 14B[44], DeepCoder 14B[51], SeedCoder 8B Reasoning[46], ’CodeV R1 Distill Qwen 7B[52]) tend to be best-performing models across benchmarks. Their particularly competitive performance is also influenced by by their larger parameter count (e.g., 685B for the two versions of DeepSeek-R1 and 235B for Qwen3-A22). However, the performance improvement on RTL design goals comes at the cost of increased computational overhead during inference, as these reasoning models produce substantially longer generations; All reasoning models were evaluated using a maximum context length of 16,384 tokens, making room for their reasoning chains, whereas the rest of models performed effectively with 2,048 tokens. Experiments to measure the impact of reasoning chain length on model performance are detailed in Appendix -E. While large reasoning models achieve the highest scores, smaller RTL-specific models emerge as a particularly efficient solution for RTL generation. LLMs with either 6B or 7B parameters, such as Origen[53] and HaVen-CodeQwen[54], demonstrate stable performance across multiple tasks, highlighting the importance of specialization for RTL tasks. This suggests two primary strategies for creating high-quality LLMs for RTL generation: either even larger general-purpose LLMs or small but highly specialized models fine-tuned using RTL data.
Figure 4: Aggregated model performance across design goals for specification-to-RTL task (VerilogEval & RTLLM). General purpose models are shown with solid lines, coding models with dashed lines, and RTL-specific models with dotted lines.
@@ III-B Model Results
According to the results of Figure 2, the single-line completion task, measured through LCA, is the most challenging of all. This might be due to the existence of long-range flow, data dependencies, and multi-file contexts [28]. The rest of design goals can be analyzed visually with a cascade plot like the one in Figure 4, which shows average model performance across STX, FNC, SYN, and PSQ for the specification-to-RTL task, evaluated across all benchmarks. Since these TuRTLe goals follow a cascading evaluation, the performance observed in this plot illustrates how RTL-generated code performance drops, as the evaluation criteria become progressively challenging. On average, models achieve a relatively high STX score of 76.36%, indicating that generating syntactically correct Verilog is not a hard challenge in most cases. When, transitioning from syntax correctness to functional correctness results suffer a substantial drop, with the FNC score averaging a 42.37% with a sharp decrease of -33.99%. This decline suggests that while models are proficient at producing Verilog code that adheres to syntax rules, they struggle to generate functionally valid implementations. These results identify which should be the immediate next goal for LLMs in the context of RTL generation, and where should efforts be put during training efforts. After FNC, it can be observed how all models perform very close to their human baseline, and no significant drops take place, neither at SYN nor at PSQ. This is likely influenced by the nature of the codes that do pass the FNC stage, which are likely to be the simplest ones, representing little challenge on the later SYN and PSQ stages (with some exceptions, see Appendix -B). Once LLMs manage to pass the FNC stage of the more challenging samples (roughly 25% of the ones currently included in TuRTLe), the SYN and PSQ goals may exhibit significant drops.
@@ IV Conclusions
This work presents TuRTLe, the first unified evaluation framework for assessing LLMs in RTL generation, running across various tasks (single-line completion, module completion, and specification-to-RTL) and design goals (line-level contextual accuracy, syntax correctness, functional correctness, synthesizability, and post-synthesis quality). TuRTLe integrates multiple benchmarks into an automated and scalable pipeline that simplifies and standardizes experimentation. The framework is extendable by design both with new benchmarks and new models. To enhance the visibility and accessibility, an online TuRTLe leaderboard is released and maintained [^4] https://huggingface.co/spaces/HPAI-BSC/TuRTLe-Leaderboard, as illustrated in Appendix -F.
Results obtained from TuRTLe-based experimentation, the largest to date on RTL generation with forty models, reveal several critical insights. Reasoning-based models achieve state-of-the-art performance, despite their higher token generation requirements and inference latency. Model type also influences task performance, as base models outperform instruction-tuned models in completion tasks, while instruct-tuned models are more effective in specification-to-RTL tasks. Small and specialized models, tuned specifically with RTL data, perform competitively even with substantially fewer parameters.
Analysis of RTL tasks indicates LLMs approach saturation on current syntax benchmarks. However, functional correctness remains challenging, highlighting what should become the short-term research priority. Our contribution extends current assessment capabilities, by introducing a novel PPA score which quantifies post-synthesis RTL quality relative to human references. The proposed evaluation scheme acts as a cascade, linking advancements across core RTL generation requirements.
Overall, this work represents a significant step towards a comprehensive evaluation of LLMs for EDA workflows, providing insights into their capabilities and limitations in RTL generation. Our findings underscore the need for more sophisticated and realistic benchmarks that better reflect real-world design challenges, including multi-module architectures and complex module interconnections. The TuRTLe framework source code is released on GitHub [^5] https://github.com/HPAI-BSC/TuRTLe, ensuring reproducibility. Future work will expand this framework and its associated leaderboard with additional benchmarks, further supporting the evaluation and development of better LLMs for hardware design automation.